OpenStudy (anonymous):

The length of a rectangle is 2 m less than twice the width. If the perimeter is 11 m, find the length and width.

OpenStudy (anonymous):

if the width is w: what's the length in terms of w?

OpenStudy (mayankdevnani):

 This is rectangle....length=2x-2, breadth=2x perimeter=11 p. of rectangle=2(l+b) =2(2x-2+2x)=11 x=1.875 l=1.75 b=3.75